What is the mass of dextrose that is in 45.6 g of a 10.3% solution of dextrose?


1
Expert's answer
2022-04-14T21:48:05-0400

w = m(solute) / m(solution)

m(dextrose) = w × m(solution) = 0.103 × 45.6 g = 4.70 g